the specimen of copper wires drawn from a large lot have the follwing breaking strength (in kg. weight): 578,572,570,568,572,578,570,596,544 determine 95% confidence interval for the true mean breaking strength
Answer #1

Ans:

1 578
2 572
3 570
4 568
5 572
6 578
7 570
8 596
9 544
mean= 572
std. dev= 13.5

n=9

df=9-1=8

critical t value=tinv(0.05,8)=2.306

95% confidence interval for mean

=572+/-2.306*13.5/sqrt(9)

=572+/-10.38

=(561.62, 582.38)